OpenStudy (loser66):

Define a non zero functional y on C^3 such that x_1= (1,1,1) and x_2= (1,1,-1), then [x1,y]=[x2,y]=0 and find all such functionals I don't know how to start Please, help

OpenStudy (loser66):

\[x_1=(\xi_1,\xi_2, \xi_3)~~and~~[x_1,y]=\xi_1-\xi_2 =0\] \[x_2=(\eta_1,\eta_2,\eta_3)~~and~~[x_2,y]=\eta_1+\eta_3=0\] then, \([x_1,y]=[x_2,y]=0\) From the method above, I can have 1/ above functional 2/\([x_1,y] =\xi_1-\xi_3\) \([x_2,y]=\eta_1-\eta_2\) 3/\([x_1,y] =\xi_2-\xi_3\) \([x_2,y]=\eta_2+\eta_3\) Am I on the right way? Pleaaaaaaaaaaaase

OpenStudy (loser66):

@amistre64 @phi

OpenStudy (amistre64):

pfft, i dont even know how to start with that either

OpenStudy (loser66):

Thanks for respond @amistre64

ganeshie8 (ganeshie8):

is this really linear algebra

OpenStudy (loser66):

abstract algebra

ganeshie8 (ganeshie8):

oh i better run lol i remember @ikram002p took this last sem

OpenStudy (loser66):

then, I am waiting for him. :)

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

whts the question

OpenStudy (loser66):

@beastonastick am I on the right way?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

if so, how to prove the linear functional is correct?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

it wont let me put the file down

OpenStudy (ikram002p):

ill work on it tomoro.. @ganeshie8 btw i took it last year spring :D

OpenStudy (loser66):

I approach the problem on another way: [x1,y]=y(1,1,1) = a1+a2+a3 =0 [x2,y]=y(1,1,-1) = a1+a2-a3 =0 solve for them, I have a3=0, a1=-a2 so, y is defined by [x,y] = a1-a2 another is [x,y]= a2-a1 @ikram002p I confused every thing. I consider dropping the class :( The most difficult thing is my prof symbolizes the concept without illustrating by numeric examples. I ....am tired with the course!!!

OpenStudy (ikram002p):

hmm it simple course oly this functional need to work on u just need to solve more of examples then ull get the idea, the rest of it is like playing with number and groups or so , dnt give up

OpenStudy (ikram002p):

ok lets try this problem using function A onto B or in ur case X onto Y 

OpenStudy (ikram002p):

\[\phi:X \rightarrow Y\] is onto if for each y in Y there is at least one x in X such that \[\phi (x)=y\]

OpenStudy (ikram002p):

so nw we need to know what kind of function is onto on \[C^3 \rightarrow C\] s.t y(x1)=y(x2) the rest is as u did lol